### v2

@@ 2.5 Performance Evaluation Metrics
Accuracy=True Positive+True NegativeTotal $\text{Accuracy}=\frac{\text{True Positive}+\text{True Negative}}{\text{Total Population}}$
Precision=True PositiveTrue Positive+False $\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}$
Recall=True PositiveTrue Positive+False $\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}$
[…]
Cross-validation is crucial in model development for two key reasons: it helps prevent over-fitting by assessing a model’s performance across different subsets of the data, and it ensures the model’s generalization ability, providing a reliable estimation of its effectiveness under various conditions. The model was trained and evaluated using k-fold cross-validation, a robust technique for assessing the generalization performance of the model. In each iteration of the k-fold cross-validation process, the dataset was partitioned into $k$ folds, and the model was trained on $k-1$ folds while being validated on the remaining fold. This process was repeated k times, ensuring that every fold had the opportunity to serve as the validation set. For each fold, the Receiver Operating Characteristic (ROC) curve was plotted, illustrating the trade-off between true positive rate and false positive rate at various thresholds. After completing the k-fold cross-validation, the individual ROC curves were aggregated, and the mean ROC curve was calculated and plotted. AUC (area under the curve) measures the entire two-dimensional area underneath the entire ROC curve. AUC provides an aggregate measure of performance across all possible classification thresholds. One way of interpreting AUC is as the probability that the model ranks a random positive example more highly than a random negative example. AUC ranges in value from 0 to 1. A model whose predictions are 100% wrong has an AUC of 0.0; one whose predictions are 100% correct has an AUC of 1.0.This comprehensive approach provides a more reliable estimation of the model’s performance, capturing its consistency across different subsets of the data and enhancing the overall assessment of its predictive capabilities.
@@ 3.2 Performance of ViT+SVM
We employed the Vision Transformer (ViT) models for feature extraction and subsequent evaluation through a Support Vector Machine (SVM) classifier. Primarily, we utilized the ViT-B/16 architecture, sourced from torchvision.models, and loaded pre-trained weights using torch.load from a specified path. The model was initialized and set to evaluation mode to facilitate inference-based feature extraction. To standardize the inputs, all images were resized to 224$\times$224 pixels and transformed into tensors. Data loaders were constructed for both the training and testing datasets, using a batch size of 32. Feature extraction was encapsulated in a dedicated function invoked separately for the training and testing data loaders. The resulting features, along with their corresponding labels, were saved to CSV files for downstream processing. This process was executed within a no-gradient context torch.no_grad() to optimize computational efficiency. The hyperparameters and implementation details used for ViT-based feature extraction and classification are summarized in Table 4.
@@ 3.3 Comparison of the Models
Our two proposed approaches are compared with a previously published model [18] on the same dataset. Table 7 presents a comparative evaluation of the three models for drought stress detection, reporting precision, recall, F1-score, and overall accuracy on the designated test set. The CNN-based framework demonstrates competitive performance, particularly in terms of stressed plant precision (0.9673). However, it lags behind in F1-score and overall accuracy when compared to the Vision Transformer with Transfer Learning. The ViT+SVM model shows relatively weaker performance, especially in classifying healthy plants, suggesting that the SVM integration may not fully exploit the representational capacity of ViTs for this task. The Vision Transformer with Transfer Learning outperforms both the CNN-based and ViT+SVM models across all metrics, making it the most effective model for distinguishing drought-stressed from healthy plants.
@@
CNN Based Framework [18] 0.9673 0.8869 0.9252 0.8203 0.9451 0.8788 0.9075

### v3

@@ 2.5 Performance Evaluation Metrics
$\text{Accuracy}=\frac{\text{True Positive}+\text{True Negative}}{\text{Total Population}}$
$\text{Precision}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Positive}}$
$\text{Recall}=\frac{\text{True Positive}}{\text{True Positive}+\text{False Negative}}$
$\text{F1-score}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
Cross-validation is crucial in model development for two key reasons: it helps prevent over-fitting by assessing a model’s performance across different subsets of the data, and it ensures the model’s generalization ability, providing a reliable estimation of its effectiveness under various conditions. The model was trained and evaluated using k-fold cross-validation, a robust technique for assessing the generalization performance of the model. In each iteration of the k-fold cross-validation process, the dataset was partitioned into $k$ folds, and the model was trained on k1k-1 folds while being validated on the remaining fold. This process was repeated k times, ensuring that every fold had the opportunity to serve as the validation set. For each fold, the Receiver Operating Characteristic (ROC) curve was plotted, illustrating the trade-off between true positive rate and false positive rate at various thresholds. After completing the k-fold cross-validation, the individual ROC curves were aggregated, and the mean ROC curve was calculated and plotted. AUC (area under the curve) measures the entire two-dimensional area underneath the entire ROC curve. AUC provides an aggregate measure of performance across all possible classification thresholds. One way of interpreting AUC is as the probability that the model ranks a random positive example more highly than a random negative example. AUC ranges in value from 0 to 1. A model whose predictions are 100% wrong has an AUC of 0.0; one whose predictions are 100% correct has an AUC of 1.0.This comprehensive approach provides a more reliable estimation of the model’s performance, capturing its consistency across different subsets of the data and enhancing the overall assessment of its predictive capabilities.
@@ 3.2 Performance of ViT+SVM
We employed the Vision Transformer (ViT) models for feature extraction and subsequent evaluation through a Support Vector Machine (SVM) classifier. Primarily, we utilized the ViT-B/16 architecture, sourced from torchvision.models, and loaded pre-trained weights using torch.load from a specified path. The model was initialized and set to evaluation mode to facilitate inference-based feature extraction. To standardize the inputs, all images were resized to $\times224$ pixels and transformed into tensors. Data loaders were constructed for both the training and testing datasets, using a batch size of 32. Feature extraction was encapsulated in a dedicated function invoked separately for the training and testing data loaders. The resulting features, along with their corresponding labels, were saved to CSV files for downstream processing. This process was executed within a no-gradient context torch.no_grad() to optimize computational efficiency. The hyperparameters and implementation details used for ViT-based feature extraction and classification are summarized in Table 4.
@@ 3.3 Comparison of the Models
Our two proposed approaches are compared with a previously published model [patra_explainable_2024] on the same dataset. Table 7 presents a comparative evaluation of the three models for drought stress detection, reporting precision, recall, F1-score, and overall accuracy on the designated test set. The CNN-based framework demonstrates competitive performance, particularly in terms of stressed plant precision (0.9673). However, it lags behind in F1-score and overall accuracy when compared to the Vision Transformer with Transfer Learning. The ViT+SVM model shows relatively weaker performance, especially in classifying healthy plants, suggesting that the SVM integration may not fully exploit the representational capacity of ViTs for this task. The Vision Transformer with Transfer Learning outperforms both the CNN-based and ViT+SVM models across all metrics, making it the most effective model for distinguishing drought-stressed from healthy plants.
@@
CNN Based Framework [patra_explainable_2024] 0.9673 0.8869 0.9252 0.8203 0.9451 0.8788 0.9075